i simply just dont understand how to solve these! iv looked online, in my notes and my math book and i dont understand how they come to the solution for these:
find the point on the line 2x+y=5 that is closest to the orgin.

Answer by rfer(16322)

Make an input/output table
graph the results
pick the spot nearest (0,0)
2x+y=5
2x+0=5
x=5/2
x=2.5
(2.5,0)
-------------
2*0+y=5
y=5
(0,5)
--------
You can pick the nearest point, but it looks like
the x intercept (2.5,0)